OpenStudy (anonymous):

find: f^-1(x) f(x)=^3√x-9

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x-9}\]?

OpenStudy (anonymous):

Yes That is the correct problem

OpenStudy (anonymous):

put \[y=\sqrt[3]{x-9}\] then either solve for \(x\) or switch \(x\) and \(y\) (because that is what the inverse does ) and solve fo \(y\) you get \[x=\sqrt[3]{y-9}\] cube to get \[x^3=y-9\] then add 9 to get \[x^3+9=y\] and that is your inverse \[f^{-1}(x)=x^3+9\]

OpenStudy (anonymous):

Oh Wow!